# STeP: A General and Scalable Framework for Solving Video Inverse Problems with Spatiotemporal Diffusion Priors

### v1

@@ 3.2 Spatiotemporal diffusion prior
Figure 3: Architecture of the spatiotemporal module in the proposed spatiotemporal UNet. Given a pretrained image diffusion UNet, we incorporate a zero-initialized temporal module with an ON/OFF switch into each 2D spatial module and initialize the additive weights ($\alpha$) to zero. Thus, it will add no effect at the start of fine-tuning and gradually learn by the video training data. The number of frames, height, and width are denoted by $n_{f}$, $n_{h}$, and $n_{w}$, respectively. The numbers of channels for input features ($f_{\text{in}}$) and output features ($f_{\text{out}}$) are denoted by $n_{\text{in}}$ and $n_{\text{out}}$, respectively.
Latent diffusion model with image encoder (P1).
@@
We start by training a VAE [35] using the standard $L_{1}$ reconstruction loss with a scaled KL divergence loss on an image dataset. The KL divergence scaling factor is set to much less than 1 to prevent excessive regularization of the latent space. This allows us to obtain an image encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Once they are trained, we fix their parameters and train a 2D UNet model ${\mathbf{s}}_{{\bm{\theta}}}({\mathbf{z}}_{t};\sigma_{t})$ using the standard denoising score matching loss. Despite recent progress in 3D spatiotemporal encoders and decoders [75, 72, 11], we opt for a 2D spatial encoder and decoder that processes each frame independently. This choice is due to efficiency considerations for the downstream PnP diffusion solver, where the decoder $\mathcal{D}$ is called multiple times during posterior sampling.
Spatiotemporal UNet as score function (P2).
@@
Leveraging recent advancements in video generation [66, 28], we use a spatiotemporal UNet architecture to parameterize the time-dependent video score function, i.e. ${\mathbf{s}}_{{\bm{\theta}}}({\mathbf{z}}_{t};\sigma_{t})\approx\nabla_{{\mathbf{z}}_{t}}\log p({\mathbf{z}}_{t};\sigma_{t})$. The key component in the architecture is a spatiotemporal module for 3D modeling, as illustrated in Fig. 3. Given a pretrained image diffusion UNet, we introduce a zero-initialized temporal module for each 2D spatial module. Specifically, for an input feature $f_{\text{in}}$, let $f_{\text{out}}$ be the output of the spatiotemporal module, with $f_{\text{spat}}$ and $f_{\text{temp}}$ representing the outputs of the spatial and temporal branches, respectively. These features are combined using an alpha blending mechanism:
@@
where $\alpha\in\mathbb{R}$ is a learnable parameter initialized as 00 in each spatiotemporal module. This design allows us to inherit the weights of the 2D spatial modules from the pretrained image diffusion model, significantly reducing the required training time. Additionally, by factorizing the 3D module into a 2D spatial module and a 1D temporal module, the spatiotemporal UNet only has marginal computational overhead compared to the original 2D UNet, striking a good balance between model capacity and efficiency.
Image-video joint fine-tuning (P3).
For compatibility with both image and video inputs, we introduce an ON/OFF switch signal in the spatiotemporal module. When the switch is set to OFF (indicating image input), the temporal module is disabled (or equivalently set $\alpha=0$). This ensures the output to $f_{\text{out}}=f_{\text{spat}}$ and reduces the spatiotemporal module to the original 2D spatial module, which processes each frame independently. During training, we initialize the weights of the spatial modules based on a pretrained image diffusion model and fine-tune all parameters of the spatiotemporal UNet using both image and video data. During fine-tuning, the model receives video data with probability $p_{\text{joint}}\in[0,1]$ and receives a pseudo video, where each frame is randomly sampled from an image dataset, with probability $1-p_{\text{joint}}$. The probability $p_{\text{joint}}$ is a tunable hyperparameter controlling the proportion of real video data in training. Pseudo video regularization helps the spatiotemporal UNet retain the spatial capabilities of the initialized spatial UNet. This strategy stabilizes training and prevents overfitting to the video dataset, proven effective in previous work [66].
3.3 Decoupled annealing posterior sampling
After obtaining a spatiotemporal diffusion prior, it is theoretically possible to combine it with any PnP diffusion solver. In this work, we employ the Decoupled Annealing Posterior Sampling (DAPS) framework, which is a novel framework for solving general inverse problems [81]. It is also easily compatible with latent diffusion models, making it an ideal choice for our purpose.
The core idea of DAPS is to sample the target latent posterior $p({\mathbf{z}}_{0}|{\mathbf{y}})$ by sequentially sampling $p({\mathbf{z}}_{t}|{\mathbf{y}})$ from $t=T$ to $t=0$. To do so, DAPS starts from $p({\mathbf{z}}_{T}|{\mathbf{y}})\approx\mathcal{N}({\bm{0}},\sigma_{\max}^{2}{\bm{I}})$ and sequentially draws a sample from $p({\mathbf{z}}_{t_{i-1}}|{\mathbf{y}})$ given a sample from $p({\mathbf{z}}_{t_{i}}|{\mathbf{y}})$ for $i=N,...,1$ based on a time schedule $\{t_{i}\}_{i=1}^{N}$. As shown by Proposition 1 of [81], this is possible if one can sample from:
Indeed, by accessing the gradient, this unnormalized distribution can be sampled by MCMC methods, such as Langevin Monte Carlo (LMC) [69] and Hamiltonian Monte Carlo (HMC) [3]. After obtaining ${\hat{\mathbf{z}}_{0}}\sim p({\mathbf{z}}_{0}|{\mathbf{z}}_{t_{i}},{\mathbf{y}})$, we can easily sample from $p({\mathbf{z}}_{t_{i-1}}|{\mathbf{y}})$ by sampling ${\mathbf{z}}_{t_{i-1}}\sim\mathcal{N}({\hat{\mathbf{z}}_{0}},\sigma_{t_{i-1}}^{2}{\bm{I}})$ due to Proposition 1 of [81]. The pseudocode and more technical details of the proposed algorithm are provided in Appendix A.
@@ 4 Experiments
Table 1: Results on Black Hole Imaging for a Test Dataset of 20 Videos. We report the average with the standard deviation in parentheses. We use d-PSNR and d-SSIM to refer to PSNR and SSIM computed over consecutive frames. Due to the high ill-posedness of black hole imaging, we select the best out of five i.i.d. posterior samples based on the lowest average $\tilde{\bm{\chi}}_{\text{cp}}^{2}$ and $\tilde{\bm{\chi}}_{\text{logca}}^{2}$ for each test video. We find that the proposed spatiotemporal prior significantly enhances the temporal consistency (see middle columns) and improves per-frame spatial consistency (see left columns) compared to the baselines. As a result, the sampled video reconstructions better align with the observations, as shown in the better data fitting $\tilde{\bm{\chi}}^{2}$ statistics.
Methods PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) d-PSNR ($\uparrow$) d-SSIM ($\uparrow$) FVD ($\downarrow$) $\tilde{\bm{\chi}}_{\text{cp}}^{2}$ ($\downarrow$) $\tilde{\bm{\chi}}_{\text{logca}}^{2}$ ($\downarrow$)
STeP (ours) 27.23 (3.26) 0.75 (0.12) 0.172 (0.077) 39.05 (4.26) 0.95 (0.04) 192.34 1.907 (1.422) 1.403 (0.589)
IDM+BCS 25.54 (2.44) 0.74 (0.09) 0.183 (0.051) 32.54 (3.99) 0.94 (0.02) 255.41 2.411 (0.219) 2.380 (0.767)
IDM 24.13 (2.30) 0.69 (0.10) 0.196 (0.061) 29.42 (2.30) 0.92 (0.05) 1336.23 3.483 (3.454) 2.789 (2.753)
Figure 4: Visual results of black hole imaging. Left: We display selected frames from the ground truth and sampled video reconstructions from each method, with a stride of 8 frames. STeP achieves the best visual quality, accurately recovering the ring radius, bright spot location, and overall black hole appearance. Right: We visualize temporal dynamics by averaging the delta between consecutive frames over 4, 16, and 64 frames, respectively. STeP faithfully recovers motion patterns that align with the ground truth. In contrast, IDM+BCS underestimates the degree of motion and IDM lacks a consistent motion pattern. These spatiotemporal structures are further evident in the averaged optical flow over the entire 64 frames, where the optical flow is estimated using a pretrained model from [59].
We demonstrate the effectiveness of STeP on two challenging scientific inverse problems: black hole imaging [22] (Sec. 4.2) and dynamic MRI [25] (Sec. 4.3). We also provide an ablation study on the effectiveness of the image-video joint fine-tuning technique in Sec. 4.4. We provide additional experimental results and visualizations in Appendix E.
4.1 Baselines & Metrics
Baselines.
We establish a comparison by introducing two baselines. The first baseline replaces the video diffusion prior with an image diffusion prior, which is applied independently to each frame (referred to as IDM). The second baseline leverages the batch-consistency sampling technique [37] with an image diffusion prior (referred to as IDM+BCS). While IDM treats each frame as an independent image inverse problem, IDM+BCS enforces temporal consistency implicitly via correlated noise.
Metrics.
We evaluate frame-wise similarity between generated and ground truth videos using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [68], and Learned Perceptual Image Patch Similarity (LPIPS) [82]. These metrics are computed independently for each frame and then averaged. We use the versions implemented in piq [33] with all images normalized to the range $[0,1]$. For gray-scale frames, we repeat them 3$\times$ along the channel dimension before calculating the LPIPS score.
To assess temporal consistency, we introduce d-PSNR and d-SSIM, which compute PSNR and SSIM over the delta between consecutive frames. These metrics are also averaged across all delta frames. Additionally, we compute the Fréchet Video Distance (FVD) [61] between the test dataset and all video reconstructions to measure distributional similarity.111 We use the following project to compute FVD: https://github.com/JunyaoHu/common_metrics_on_video_quality.
Finally, we report the measurement data consistency using domain-specific metrics. For dynamic MRI, we report the mean squared error $\|\mathcal{A}({\mathbf{x}})-{\mathbf{y}}\|_{2}$ as data misfit. For black hole imaging, we use the $\bm{\chi}^{2}$ statistic (referred to Eq. 16 for detailed definition) on two closure quantities: the closure phase ($\bm{\chi}_{\text{cp}}^{2}$) and log closure amplitude ($\bm{\chi}_{\text{logca}}^{2}$). A $\bm{\chi}^{2}$ value close to 1 indicates good data fitting (refer to Appendix B for more detail). To facilitate a comparison between underfitting $(\bm{\chi}^{2}>1)$ and overfitting $(\bm{\chi}^{2}<1)$, we report a unified metric defined as:
Table 2: Results on Dynamic MRI with 6$\times$ acceleration for a Test Dataset of 20 Videos. We report average PSNR, SSIM of the real and imaginary components and do similarly for d-PSNR, d-SSIM. The standard deviations are included in parentheses. LPIPS and FVD scores are calculated over the complex amplitude. The results show that by leveraging the proposed spatiotemporal prior, STeP consistently improves both temporal and per-frame spatial consistency.

### v2

@@ 3.3 Efficient training of spatiotemporal diffusion priors
Figure 3: Architecture of the spatiotemporal module. Given a pretrained image diffusion UNet, we add a zero-initialized temporal module with an ON/OFF switch to each 2D spatial module and initialize the additive weight $\alpha$ to zero. Thus, it will add no effect at the start of fine-tuning and gradually learn by the video training data. The number of frames, height, and width are denoted by $n_{f}$, $n_{h}$, and $n_{w}$, respectively. The numbers of channels for input features ($f_{\text{in}}$) and output features ($f_{\text{out}}$) are denoted by $n_{\text{in}}$ and $n_{\text{out}}$, respectively.
Step 1: Training a latent diffusion model (LDM) as an image prior
@@
We start by training a VAE (kingma2022autoencodingvariationalbayes, ) using the standard $L_{1}$ reconstruction loss with a scaled KL divergence loss on an image dataset. The KL divergence scaling factor is set to much less than 1 to prevent excessive regularization of the latent space. This allows us to obtain an image encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Once they are trained, we fix their parameters and train a 2D UNet model ${\mathbf{s}}_{{\bm{\theta}}}({\mathbf{z}}_{t};\sigma_{t})$ using the standard denoising score matching loss vincent2011aconnection . Despite recent progress in 3D spatiotemporal encoders and decoders (yang2024cogvideoxtexttovideodiffusionmodels, ; wu2024improvedvideovaelatent, ; chen2024odvaeomnidimensionalvideocompressor, ), we opt for a 2D spatial encoder and decoder to process each frame independently. This choice is due to efficiency considerations, as the decoder $\mathcal{D}$ will be called repeatedly in DAPS during inference.
Step 2: Turning an image prior into a spatiotemporal prior
@@
Upon obtaining an LDM as an image prior, we use a spatiotemporal UNet architecture to parameterize the time-dependent video score function, i.e. ${\mathbf{s}}_{{\bm{\theta}}}({\mathbf{z}}_{t};\sigma_{t})\approx\nabla_{{\mathbf{z}}_{t}}\log p({\mathbf{z}}_{t};\sigma_{t})$, leveraging recent advancements in video generation (wang2023modelscopetexttovideotechnicalreport, ; ho2022videodiffusionmodels, ). The key component in the architecture is the spatiotemporal module for 3D modeling, as illustrated in Figure 3. Given a pretrained IDM with a 2D UNet, we introduce a zero-initialized temporal module for each 2D spatial module within the UNet. Specifically, for an input feature $f_{\text{in}}$, let $f_{\text{out}}$ be the output of the spatiotemporal module, with $f_{\text{spat}}$ and $f_{\text{temp}}$ representing the outputs of the spatial and temporal branches, respectively. These features are combined using an $\alpha$-blending mechanism:
@@
where $\alpha\in\mathbb{R}$ is a learnable parameter initialized as 00 in each spatiotemporal module. This design allows us to inherit the weights of the 2D spatial modules from the pretrained IDM, significantly reducing training time. Additionally, by factorizing the 3D module into a 2D spatial module and a 1D temporal module, the spatiotemporal UNet has marginal computational overhead compared to the original 2D UNet, striking a good balance between model capacity and efficiency.
Step 3: Image-video joint fine-tuning
To further improve performance and be compatibile with both image and video inputs, we introduce an ON/OFF switch signal in the spatiotemporal module. When the switch is set to OFF (indicating image inputs), the temporal module is disabled (or equivalently $\alpha=0$). This ensures that $f_{\text{out}}=f_{\text{spat}}$ and reduces the spatiotemporal module to the original 2D spatial module, which processes each frame independently. During training, we initialize the weights of the spatial modules based on a pretrained image diffusion model and fine-tune all parameters of the spatiotemporal UNet using both image and video data. During fine-tuning, the model receives video data with probability $p_{\text{joint}}\in[0,1]$ and receives images (with switch set to OFF accordingly) with probability $1-p_{\text{joint}}$. The probability $p_{\text{joint}}$ is a tunable hyperparameter controlling the proportion of real video data in training. Pseudo video regularization helps the spatiotemporal UNet retain the spatial capabilities of the initialized spatial UNet. This strategy, proven effective in previous work (wang2023modelscopetexttovideotechnicalreport, ), stabilizes training and prevents overfitting to the video dataset.
@@ 4 Experiments
4.1 Tasks and setup
We consider two scientific video inverse problems: black hole video reconstruction akiyama2022firstIII and dynamic MRI gamper2008compressed . Although both are scientific imaging tasks, they have significantly different characteristics. Black hole video reconstruction involves simple spatial structures (usually a ring structure) but complex temporal dynamics that obey physical constraints. In contrast, dynamic MRI requires higher spatial fidelity with relatively simpler temporal dynamics, such as periodic heartbeat motion.
Black hole video reconstruction
We consider the problem of observing the Sagittarius A* black hole using the Event Horizon Telescope (EHT) array in 2017 akiyama2022firstIII . The measurements consist of sparse vectors with a dimensionality of 1,856 derived from an underlying black hole video comprising 64 frames, each with a spatial resolution of 256$\times$256 pixels, capturing the black hole’s dynamics during the 100-minute observation period. These measurements are given by calculating the closure quantities based on the complex visibilities (a detailed description of the problem is available in Appendix B.1). For the training dataset, we employ the general relativistic magnetohydrodynamic (GRMHD) simulations (wong2022patoka, ) of Sagittarius A*, covering various black hole models and observational conditions. The dataset comprises 648 simulated black hole videos and 50,000 black hole images.
Dynamic MRI
We consider a standard compressed sensing MRI setup with two acceleration scenarios: 8$\times$ acceleration with 12 auto-calibration signal (ACS) lines and 6$\times$ acceleration with 24 ACS lines. Further details are provided in Appendix B.2. We utilize the publicly available cardiac cine MRI dataset from the CMRxRecon Challenge 2023 wang2024cmrxrecon for training. This dataset includes 3,324 cardiac MRI sequences, each containing fully sampled, ECG-triggered $k$-space data from 300 patients, featuring various canonical cardiac imaging views. Each sequence is processed into a video consisting of 12 frames with a spatial resolution of 192$\times$192 pixels. We construct the image dataset by extracting the individual frames from the videos, resulting in 39,888 images in total.
4.2 Baselines and our method
Recall from Sec. 3 that our method leverages a learned spatiotemporal prior to avoid explicitly estimating temporal dynamics from measurements at inference time. To evaluate the effectiveness of this strategy, we compare our method to existing IDM-based approaches. Based on how temporal information is incorporated, we categorize these baseline methods into two groups as follows.
Group 1: simple heuristics
We consider two baselines: BIS kwon2025solving and BCS kwon2025solving . Batch independent sampling (BIS) reconstructs each frame independently, while batch consistent sampling (BCS) implicitly incorporates a static temporal prior to enforce consistency across frames.
Group 2: noise warping
We include another line of baselines: $\int$-noise chang2024how ; deng2025infiniteresolution , and GP-Warp daras2024warped . These baselines enforce temporal consistency by warping the noise using optical flow estimated directly from the measurements. Besides, we include more conventional warping strategies such as Bilinear, Bicubic, Nearest from work chang2024how . For the black hole video reconstruction task, a direct inversion from the sparse measurement vector is infeasible. Thus, to demonstrate the upper-bound performance of these methods, we leverage the ground truth video as an oracle to derive the optical flow. For the dynamic MRI task, we utilize a naive inversion obtained via inverse Fourier transformation to estimate the optical flow. We follow the methodology of chang2024how , leveraging a pretrained model teed2020raftrecurrentallpairsfield to extract optical flow.
Ours: two variants of STeP
We evaluate two variants differentiated by their spatiotemporal priors: STeP (video-only), which uses a spatiotemporal prior trained exclusively on video data, and STeP (image-video joint), which initializes with a pre-trained IDM and undergoes joint image-video fine-tuning. We follow the Sec. C to train our priors until convergence. The detailed training hyperparameters are summarized in Table 3. To make a fair comparison, we run all the baselines with DAPS zhang2024improving . The detailed implementation of each baseline can be found at Appendix A and B.3.
4.3 Metrics
We evaluate our results on three aspects: (1) spatial similarity, (2) temporal consistency, and (3) measurement data fit.
Spatial
Spatial similarity is assessed by calculating Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) (wang2004image, ), and Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018lpips, ). Metrics are computed per frame and averaged across all frames. We utilize implementations from piq (kastryulin2022piq, ), normalizing images to the range $[0,1]$. For grayscale images, frames are replicated across three channels before computing LPIPS scores.
Temporal
Temporal consistency is quantified using delta-based PSNR (d-PSNR) and delta-based SSIM (d-SSIM), measuring the similarity of normalized differences between consecutive frames, with results averaged over all frames. Additionally, we compute the Fréchet Video Distance (FVD) (unterthiner2018towards, ) between the reconstructed videos and our test dataset to evaluate distributional similarity.
Data fit
Lastly, we report measurement data fit using task-specific metrics. For the black hole video reconstruction task, we employ the unified average $\bm{\tilde{\chi}}^{2}$ statistic (defined in Equation (13) in Appendix), where values closer to 1 indicate better data fidelity. For dynamic MRI, we measure data misfit by computing the mean squared error in measurement space.
Table 1: Quantitative results on black hole video reconstruction and dynamic MRI. We compare our method against baselines by reporting the mean and standard deviation (shown in parentheses) of selected evaluation metrics computed over 10 test videos (FVD is reported without standard deviation since it evaluates the set of 10 videos collectively). The results clearly demonstrate that by leveraging the spatiotemporal prior, STeP consistently achieves improvements in both spatial quality and temporal consistency relative to baseline methods.